INTERNAL ROTATION IN CF_{3}I⋯NH_{3} AND CF_{3}I⋯N(CH_{3})_{3} PROBED BY CP-FTMW SPECTROSCOPY

\underline{N. R. WALKER}, S. L. STEPHENS AND A. C. LEGON, School of Chemistry, University of Bristol, Bristol, BS8 1TS, U.K..

The pure rotational spectra of CF_{3}I⋯NH_{3} and CF_{3}I⋯N(CH_{3})_{3} have been measured by chirped-pulse, Fourier transform microwave (CP-FTMW) spectroscopy between 7 and 18.5 GHz. Both molecules are generated by supersonic expansion of a gas sample containing a small percentage of each precursor in a balance of argon. The spectra of both complexes are consistent with C_{3v} prolate symmetric top structures. The observed spectrum of CF_{3}I⋯NH_{3} displays evidence for internal rotation of NH_{3} about the principal axis. More than one hundred transitions of CF_{3}I⋯NH_{3} have been assigned to the internal rotor A state allowing rotational, centrifugal distortion constants and a nuclear quadrupole coupling constant for the iodine atom to be determined for this state. Measurements performed using a Balle-Flygare FTMW spectrometer further allow determination of a nuclear quadrupole coupling constant for the ^{14}N nucleus. Many transitions in the spectrum of the CF_{3}I⋯^{15}NH_{3} isotopologue have also been measured and the length of the halogen bond between the iodine and nitrogen atoms has been determined. Measurements of hyperfine components in nine different J′′←J′ transitions of CF_{3}I⋯N(CH_{3})_{3} have allowed assignment of the spectrum of this complex to determine rotational, centrifugal distortion and nuclear quadrupole coupling constants.